HETAV PATEL

Grade 11,

If Secent theta+tan theta=1/5 then sin theta is equal to dash

If Secent theta+tan theta=1/5 then sin theta is equal to dash

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
It becomes on simplifying that
(1 + sin@)/ cos@= 1/5
Now let  sin@ = x
Hence
 
5(1+x) = \sqrt(1 -x²)
On squaring
 
25(1 +x)² = 1 -x²
25 + 25x² + 50x= 1 -x²
26x² + 50x  +24 = 0
26 x² + 26x + 24 x +24 = 0
(26x +24) (x+1) = 0
x = -1 (not possible)
And x = -24/26
 
Regards
Arun (askIITians forum expert)